OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a yellow marble, do not replace it, and pick a red marble?

OpenStudy (anonymous):

1/15?

OpenStudy (anonymous):

sorry i was getting something

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first pick: 3 yellow / 10 total = 3/10 then for the second pick, 2 yellow / 9 total = 2/9 combined probability is (3/10) / (2/9) = 6/90 = 1/15 So the answer is 1/15

OpenStudy (anonymous):

I have one last qeustion What is the distance between P(0, 14) and Q(5, -2)?

OpenStudy (anonymous):

post this different question...